how many amperes of current should be passed when aqueous solution of Na2SO4 is electrolysed between graphite electrodes at 300 k temperature so that 250ml min^-1 O2 gas is obtained under 1.0 bar pressure??

Grade:11

1 Answers

Rituraj Tiwari
3 years ago
Convert the initial conditions to standard conditions (temperature and pressure)

P1V1/T1 = P2V2/T2

P1 = 1 bar = 750mmHg

V1 = 250ml = 250 cm^3

T1= 300K

P2 = 760mmHg

V2 = ?

T2 = 273 K

(750x250)/300 = (760xV2)/273

V2 = (750X250X273)/(300X760)

V2 = 224.51 cm^3

1 mole of gas occupies 22400cm^3 at STP

Moles of Oxygen = 224.51/22400 = 0.001 moles

One mole of oxygen is discharged by 4 moles of electrons

Moles of electrons = 4 x 0.001 = 0.004

Convert moles to Coulombs

1 moles of electrons = 96485 Coulombs

Number of Coulombs = 0.004 x 96484 = 386.8C

Coulombs (Q) = Current (I) x Time (t)

Time = 1 minute = 60 seconds

Current (Amps) = Q/t = 386.8/60

Current = 6.44 Amps